# The $\infty$-S test via regression quantile affine LASSO

### v1

@@ 5.3 The total variation $\infty$-Sτtest
Figure 2: Left plot: power of the total variation $\infty$-Sτ test and $\infty$-rankSτ test for $\tau=0.8$ as a function of the amount of the heteroschedasticity index $\delta\in[0,2]$. The horizontal dotted line is the nominal level $\alpha=0.05$. Right plot: example of data for $\delta=2.0$ where the test rejects $H_{0}$. The vertical dotted line is the estimated change location.
We consider a signal $\boldsymbol{y}=(y_{1},\ldots,y_{n})$ from heteroschedastic independent realizations centered around zero of length $n=100$ and with standard deviation $\sigma$ for $i=1,\ldots,n/2$ and $\sigma+\delta$ for $i=n/2+1,\ldots,n$. Considering the upper quantile with $\tau=0.8$, a quantile change occurs at the midpoint of the data when $\delta>0$ and the larger $\delta$, the bigger the heteroschedasticity (that is, the further away from $H_{0}:$ data are homoschedastic). We simulate errors $\boldsymbol{e}$ as realization from a Student distribution with 10 degrees of freedom rescaled to have unit standard deviation. We use the $\infty$-Sτ test (and its rank version) with $\tau=0.8$ to test at the level $\alpha=0.05$ the null hypothesis of constant $\tau$-quantile. The left plot of Figure 2 reports the power of the two tests as a function of $\delta\in[0,2]$. We observe that both tests have the nominal level $\alpha=0.05$ at $\delta=0$. The $\infty$-rankSτ test is slightly more powerful than the $\infty$-Sτ test. The left plot of Figure 2 shows data for $\delta=2.0$ for which the $\infty$-Sτ test rejects $H_{0}$, and the vertical line visualizes the estimated change location, namely the index of the largest value in $|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}^{\tau}|$. On this high-dimensional problem ($n=p=502$ and $m=501$), we were not able to run the standard sign test of Gutenbrunner et al. (1993) with the quantreg library (Koenker, 2024).
We apply the quantile total variation $\infty$-Sτ test to analyse the time series of the amazon daily log-returns. Focusing on the lower tail which corresponds to drops in the AMZN stock value, we test for constant $\tau$-quantile for $\tau=0.1$. The $\infty$-Sτ test is rejected with an estimated p-value of zero. The top left plot of Figure 3 shows the histogram estimate of the density $f_{S^{\tau}}$ of the test statistic and the critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$. Below this plot, Figure 3 shows the regression quantile affine LASSO path, along with the critical value $c_{\alpha}^{\tau}$ and the zero-thresholding function $\lambda_{0}^{\tau}(\boldsymbol{y},X,A)$ evaluated at the data $\boldsymbol{y}$. The right plots of Figure 3 show the raw AMZN time series data (top) and the estimated regression quantile affine LASSO estimate (bottom).
@@ 5.4 Robustness of the level to non-Gaussian errors
Under the null $H_{0}:A{\boldsymbol{\beta}}={\boldsymbol{b}}$ and for a nominal level $\alpha=0.01$, we show the effective level of the $F$-test and the $\infty$-S test as a function of the degrees of freedom for Student errors d.f. $\in\{1,2,3,4,5,10,100\}$. We choose $n=100$, $p=20$, $A$ is the first $m=5$ rows of the finite difference matrix of Section 5.3 and ${\boldsymbol{\beta}}=(3,3,4,4,\boldsymbol{h})$, where $\boldsymbol{h}$ are $16$ samples from ${\rm N}(0,5^{2})$; so $A{\boldsymbol{\beta}}$ has two null entries. In a robustness analysis of the F-test, Ali and Sharma (1996) observe that major determinant of the sensitivity to nonnormality of the errors is the extent of the nonnormality of the regressors or the extent of presence of ‘leveraged’ (influential) observations. So we simulate the entries of the matrix $X$ as a sample from $T$, where $T$ follows a student distribution with 2 degrees of freedom; $X$ is then standardized by $\sqrt{n}$.
@@
Over-shooting the nominal level leads to too many type I errors, hence rejecting too many null hypotheses, that is, making too many false discoveries.
@@ 6 Conclusions
The $\infty$-S test allows to test general linear null hypotheses for linear models. The formula is simple to implement and can handle high-dimensional problems. We derived also the test for testing quantiles and shown that other norms than the $\ell_{1}$-norm could be employed, leading to other tests.
The $\infty$-S tests can be used to derive confidence regions that are dual to tests. The $(1-\alpha)$-confidence region for $A\boldsymbol{\beta}$ based on the $\infty$-S test consists of all vectors $\boldsymbol{b}$ that are not rejected for $H_{0}:A\boldsymbol{\beta}=\boldsymbol{b}$. In particular, to test $H_{0}:\ \beta_{j}=b_{j}$ for some fixed $j\in\{1,\ldots,p\}$, one uses $A=\boldsymbol{e}_{j}^{\rm T}$, the $j$th canonical vector, and the calculated statistic is $s=|\boldsymbol{x}_{j}^{\rm T}\boldsymbol{\omega}|$, where $\boldsymbol{\omega}={\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j})$. So the $(1-\alpha)$-confidence interval for $\beta_{j}$ is the set of all $b_{j}$ such that $|\boldsymbol{x}_{j}^{\rm T}{\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j}(b_{j}))|\leq c_{\alpha}$, where $\hat{\boldsymbol{\beta}}_{-j}(b_{j})\in\arg\min_{\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p-1}}\|\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\tilde{{\boldsymbol{\beta}}}\|_{1}$ and $X_{-j}$ is the $X$ data matrix without its $j$th column.
The $\infty$-S test is available in the qut package in R.
@@ 7 Acknowledgement
This work was partially supported by Swiss National Science Foundation and the Academic Rising Star Project No. 2024-YJRC-LXY-005 of the National University of Defense and Technology, China.

### v2

@@ 5.4 The total variation $\infty$-Sτtest
We apply the quantile total variation $\infty$-Sτ test to analyse the time series of the amazon daily log-returns. Focusing on the lower tail which corresponds to drops in the AMZN stock value, we test for constant $\tau$-quantile for $\tau=0.1$. The $\infty$-Sτ test is rejected with an estimated p-value of $0.0132$. The top left plot of Figure 3 shows the histogram estimate of the density $f_{S^{\tau}}$ of the test statistic (by Monte Carlo sampling of the test statistic $S^{\tau}$ under the null hypothesis) and the critical value $c_{\alpha}^{\tau}$ for $\alpha=0.05$. Right below this plot, Figure 3 shows the quantile affine LASSO path, along with the critical value $c_{\alpha}^{\tau}$ and the zero-thresholding function $\lambda_{0}^{\tau}(\boldsymbol{y},X,A)$ evaluated at the data $\boldsymbol{y}$. The right plots of Figure 3 show the raw AMZN time series data (top) and the quantile affine LASSO estimate for $\lambda=c_{\alpha}^{\tau}$ (bottom), which corresponds to the quantile universal threshold estimate (Giacoetal17) of the $\tau$-quantile. The single jump points to a potential change of regime moving to a less volatile return.
@@ 5.5 Robustness of the level to non-Gaussian errors
Under the null $H_{0}:A{\boldsymbol{\beta}}={\boldsymbol{b}}$ and for a nominal level $\alpha=0.01$, we show the effective level of the $F$-test and the $\infty$-S test as a function of the degrees of freedom for Student errors d.f. $\in\{1,2,3,4,5,10,100\}$. We choose $n=100$, $p=20$, $A$ is the first $m=5$ rows of the finite difference matrix of Section 5.4 and ${\boldsymbol{\beta}}$ are $p$ samples from the standard Gaussian distribution; $\boldsymbol{b}$ is calculated as $A{\boldsymbol{\beta}}$. In a robustness analysis of the F-test, Ali and Sharma (1996) observe that major determinant of the sensitivity to nonnormality of the errors is the extent of the nonnormality of the regressors or the extent of presence of ‘leveraged’ (influential) observations. So we simulate the entries of the matrix $X$ as a sample from $T$, where $T$ follows a student distribution with 2 degrees of freedom.
@@
Over-shooting the nominal level leads to too many type I errors, hence rejecting too many null hypotheses, that is, making too many false discoveries. Here the two $\infty$-S tests match the nominal level, regardless of the degree of freedom of the Student errors. For this simulation, the implementation of the quantreg package does not return a p-value for $\chi^{2}$ sign test.
@@ 6 Conclusions
The $\infty$-Sτ test allows to test general linear null hypotheses for linear models and for any quantile $\tau$. The test function is simple to implement and can handle high-dimensional problems. We emphasize the $\infty$-S test which corresponds to the $\ell_{1}$-LASSO penalty, and showed that other norms than the $\ell_{1}$-norm could be employed, leading to other tests. Sardy et al. (2022) show that using $\ell_{1}$ leads to high power under sparse alternatives (that is, most entries, but not all, of $A\boldsymbol{\beta}-\boldsymbol{b}$ are zero when $H_{1}$ is true), and, under dense alternatives (that is, most or all entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ are non-zero), that the sign test based on $\ell_{2}$ has more power than the $\infty$-S test. To get nearly the best power with a single test regardless whether the alternative hypothesis is sparse or dense, Sardy et al. (2022) propose the $\oplus$-test.
The $\infty$-S test can be used to derive confidence regions that are dual to tests. The $(1-\alpha)$-confidence region for $A\boldsymbol{\beta}$ based on the $\infty$-S test consists of all vectors $\boldsymbol{b}$ that are not rejected for $H_{0}:A\boldsymbol{\beta}=\boldsymbol{b}$. In particular, to test $H_{0}:\ \beta_{j}=b_{j}$ for some fixed $j\in\{1,\ldots,p\}$, one uses $A=\boldsymbol{e}_{j}^{\rm T}$, the $j$th canonical vector, and the calculated statistic is $s=|\boldsymbol{x}_{j}^{\rm T}\boldsymbol{\omega}|$, where $\boldsymbol{\omega}={\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j})$. So the $(1-\alpha)$-confidence interval for $\beta_{j}$ is the set of all $b_{j}$ such that $|\boldsymbol{x}_{j}^{\rm T}{\rm sign}(\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\hat{\boldsymbol{\beta}}_{-j}(b_{j}))|\leq c_{\alpha}$, where $\hat{\boldsymbol{\beta}}_{-j}(b_{j})\in\arg\min_{\tilde{\boldsymbol{\beta}}\in\mathbb{R}^{p-1}}\|\boldsymbol{y}-\boldsymbol{x}_{j}b_{j}-X_{-j}\tilde{{\boldsymbol{\beta}}}\|_{1}$ and $X_{-j}$ is the $X$ data matrix without its $j$th column.
The $\infty$-S test is available and the research is made reproducible in the Stest package in R, which can be downloaded from https://github.com/StatisticsL/Stest.
@@ 7 Acknowledgement
This work was partially supported by the Swiss National Science Foundation grant 200021E_213166 and the Academic Rising Star Project No. 2024-YJRC-LXY-005 of the National University of Defense and Technology, China.